A Complete Bibliography of Publications in Internet Histories

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

26 October 2023 Version 1.01

Title word cross-reference

$27.95 [169].

/dev/null [15].

1916 [48]. 1950s [104]. 1980s [104]. 1990 [91]. 1990s [72, 115, 53].

2008 [149]. 2017 [60]. 2018 [89]. 25 [62]. 2channel [146].

3.0 [71].

4 [70].

5 [42]. 5-Star [42].

6 [179].

9 [189]. 90s [51]. 978 [189]. 978-1-5095-4108-9 [189].

Aaron [196]. abandoned [94]. Abidin [189]. abundance [102]. academia [194]. academic [98, 180]. access [175]. accordingly [16]. action [110]. activism [149]. adoption [186]. African [14]. Against [133]. Age [111, 149, 28, 71, 92, 102, 135, 112]. agenda [19]. Alan [49]. Aleksandra [129]. Aleph [171]. Aleph-bet [171]. Alexandre [67]. algorithmic [24, 100]. alt [145]. alt-right [145]. alternative [55]. am [138]. ambiguity [40]. America [32, 52]. American [126]. Amsterdam [16].

Analysing [132]. analysis [107, 191]. analytics [128]. Anarchy [199]. Andrew [61]. Anglophone [56]. anime [177]. Ann [189]. années [91]. anonymity [164]. antecedents [119]. anxieties [166]. approach [84, 182]. approaches [134]. archaeological [182]. Archaeology [16, 76]. archival [134]. archive [200, 48, 46, 47, 160, 17]. archived [181, 92]. Archives [179, 187, 39, 167, 83]. archiving [84, 181, 180, 160]. Armed [46]. ARPANET [44, 68, 65, 66, 69, 67, 64, 151]. arrangements [105]. art [81]. artificial [71]. Arwid [50]. Asian [141, 151]. assets [108]. asymmetrical [11]. author [188]. authoritarianism [133]. authority [178]. authority-building [178]. Automatic [168]. autopsy [164]. award [113, 176]. awful [191]. awry [164].

Back [67, 164]. band [77]. bans [165, 159]. Baran [27]. Barlow [172]. based [134, 119, 13]. BBS [149]. be [16, 160]. became [65]. Becoming [43]. before [105]. behavior [21]. Behind [57]. behold [33]. Beijing [149]. being [24, 71]. Benjamin [34]. Bernhard [100]. bet [171]. Between [184, 144]. Beyssac [57]. bias [131]. biopolitics [159]. birth [200, 55]. Bit [61]. Blacksburg [162]. blasphemy [3]. Blogging [70]. Bodies [118]. Book [21, 49, 60, 91, 189, 28, 169, 89, 70, 88, 123, 90, 174, 35, 22, 34]. Bory [123]. both [54]. boundary [66]. Brazil [116]. Breaking [18, 198]. breakup [194]. Broad [77]. Brügger [90]. Building [37, 55, 178]. Buzz [187].

C. [21]. Cambridge [169, 89, 174]. Can [5, 10, 160]. Capitalism [50, 111]. Capitol [179]. CAPTCHA [127]. Capturing [48]. career [113, 176]. Cascading [186]. case [193, 23, 40, 13]. cave [7]. Central [32, 52]. CERN [200]. Ceruzzi [89]. challenges [108, 13]. changing [95, 117, 158, 185]. character [13, 171]. character-based [13]. Chinese [149, 147, 150, 152]. Choking [148]. Chon [151]. Christian [56]. circle [150]. city [16, 182]. civil [114]. class [119]. class-based [119]. Classification [168]. closed [95]. Club [138]. co [41, 188]. co-author [188]. co-production [41]. code [85, 171]. coll [91]. collaboratory [31]. collection [200]. comes [135]. commemoration [48]. comment [185]. commentary [128]. commercial [117]. commercialized [115]. common [172, 4]. communication [117, 120, 180, 135, 30]. Communications [88, 107]. communities [41, 185, 184]. community [145]. comparison [75]. competition [104]. Computational [80, 79, 78, 83]. Computer [138, 194, 136, 52, 95]. computers [120]. computing [95, 7, 142]. confidential [21]. connected [33]. connecting [52]. consecration [172]. construction [68, 47, 200, 91]. contemporary [192]. Content [126, 125]. Controverses [91]. controversy [99]. Conversation [44, 68, 81, 69, 151]. Conversations [27]. conversion [26]. Cookies [99, 105]. copyright [126]. Corporation [139]. Correction [140]. Couchsurfing [158]. counterculture [58]. craigslist [130, 75]. creation [178]. Creative [152]. criminologist [38]. cross [84]. cross-platform [84]. Crypto [199, 29]. crypto-discourse [29]. Cryptoeconomics [199]. Crystal [189]. CSS [186]. cultural [74, 160, 5]. culture [196, 54]. Cultures [101, 81, 146, 58]. current [195]. curve [122]. cyberactivism [116]. cybercrime [38]. cyberlibertarian [172]. cyberlibertarianism [196]. cyberspace [56]. cynical [146]. Cypherpunk [173, 133].

Daemons [88]. dahs [171]. Danescu [110]. Danish [132]. dans [91]. Dark [87]. data [16, 131]. Dating [183, 182]. days [44]. Dead [155, 154, 166]. Dead-and-dying [155, 154]. death [156, 159, 161, 167]. deaths [158]. debate [37]. debates [185]. decade [51]. decline [154]. Defence [46]. defending [55]. define [29]. defining [2]. democracy [42]. Demonstrating [186, 18]. describing [162]. design [68, 8, 186]. designed [131]. destruction [167]. dev [15]. Development [137, 11, 200, 73, 157, 46, 30]. developmental [144]. devices [66]. Diagnosing [156]. dialectics [195]. Digital [101, 146, 88, 16, 108, 95, 19, 147, 182, 154, 25, 92, 55, 12, 58, 197, 112, 152, 21]. d’Internet [91]. direct [42]. directed [33]. discontinued [94]. discourse [95, 29, 42]. discourses [177]. Discursive [149]. discussion [79, 109]. disembodiment [118]. Disqus [185]. dissolution [117]. dits [171]. dits-and-dahs [171]. divides [12]. Do [177]. doc [79, 109]. documentaries [17]. documenting [179]. does [156]. Doing [17, 92, 157]. Don’t [164, 118]. Driscoll [60]. during [149]. Dutch [76]. dying [155, 154]. dynamic [16, 19]. dynamics [178].

Early [13, 44, 113, 176, 118, 134, 144, 180, 119, 54, 56]. East [11]. Easter [48]. EC [104]. ecology [158]. economic [105]. economy [175]. ecosystems [184]. edited [169, 90]. Editorial [103, 45]. efficacy [114]. Electronic [162, 143]. Elena [110]. Elephants [80]. Élie [69]. emergence [106]. Emigration [192]. encrypted [85]. encryption [29]. end [156, 85, 157, 122]. end-to-end [85]. engineer [38, 138]. engines [24]. epistemic [128]. era [142]. Ervin [61]. essay [119]. ethics [56]. Études [91]. EU [103, 96]. eulogies [167]. Euronet [96, 76]. Europe [108]. European [107, 106, 109]. evaluating [134]. Everything [160]. Evolution [86, 31, 132, 73, 152]. examining [119]. exclusionary [177]. exes [161]. exist [177]. Experimenting [83]. extremist [146].

Fabrique [91]. face [95]. Facebook [73]. facto} [85]. facts [47]. failure [154]. fall [72, 163]. Fandom [101]. fandoms [159]. fans [53, 177]. features [198]. feeling [9]. Felsenstein [138]. female [53, 177]. Fenwick [88]. fervent [145]. Fi [174]. field [110, 120, 135]. Fight [111]. Film [33]. filtering [126]. financialisation [10]. first [194, 62]. flame [39]. Football [101]. forced [26]. Forces [46]. Forensic [134]. Forgotten [161, 37, 142, 94]. found [84]. foundation [115]. four [158]. fragility [35]. frameworks [181]. francaise [91]. France [55, 70]. fraught [119]. free [196, 138]. Freedom [50, 196, 29]. Freenet [87]. French [121, 187, 69]. Friendster [156]. front [157]. front-end [157]. Frontier [111]. Future [139, 111, 40, 42].

G [49]. Games [61]. Garden [186]. Geert [58]. gender [166]. gendered [119]. genealogical [38]. Genealogies [125]. Genealogy [200]. genesis [192]. GeoCities [167]. Gérard [69]. global [114, 97]. globalizing [148]. godfather [151]. gone [164, 161]. Goodbye [121]. Google [137]. governance [115, 114, 199]. GPS [89]. greatest [65]. Gross [49]. Guy [32].

hacking [195]. Hagiography [3]. Han [13, 22]. handbook [90]. Hangul [144]. Hargittai [21]. Harmon [49]. harvest [76]. Hauben [188]. Having [54]. healthy [185]. Hebrew [171]. Hendry [189]. her [110]. Herding [80]. Herzog [33]. hidden [53]. historians [39]. historical [47, 191, 102, 43, 100]. histories [2, 79, 78, 176, 6, 10, 141, 3, 94, 14, 18, 62, 8, 113, 1]. History [102, 133, 19, 190, 29, 59, 134, 53, 82, 198, 7, 127, 96, 97, 144, 37, 25, 188, 92, 199, 116, 170, 12, 17, 67, 5, 112, 34, 178, 90].

Homebrew [138]. homogenization [170]. hope [175]. hopes [72]. Hot [65]. Hourtin [121]. HTML [98]. Human [111, 24, 136, 71]. Humanities [49]. hybrid [31].

I2P [87]. Ian [90]. idea [104]. ideas [185]. identification [125]. Identities [101]. identity [105]. ideological [24]. Ideologies [123]. ideology [172, 173]. If [139, 5]. Ilbe} [145]. Imaginary [123, 148]. Imagining [180, 157]. impact [188, 177]. important [23]. IMPs [66]. Ina [91]. Incomplete [169]. incunables [76]. indentured [166]. India [148, 197]. inequality [175]. influence [181]. influences [173]. Infonie [26]. information [196, 110]. infrastructural [40, 154]. infrastructure [193, 132, 181, 73]. infrastructures [66]. inherit [97]. Innovation [122]. Instagram [43]. instrumentalised [136]. intangible [108]. intellectual [133, 199]. intelligence [71]. inter [96, 66]. inter-structuralism [66]. interfaces [24]. Interfacing [58]. international [115]. Internet [49, 60, 123, 2, 11, 193, 23, 121, 47, 130, 79, 78, 176, 196, 40, 10, 103, 86, 132, 19, 106, 141, 190, 115, 114, 29, 77, 134, 53, 82, 7, 120, 96, 97, 26, 144, 180, 37, 119, 13, 192, 88, 42, 160, 151, 148, 55, 135, 17, 3, 51, 67, 109, 32, 52, 143, 94, 9, 197, 5, 14, 18, 150, 152, 30, 34, 8, 1, 113, 188]. Interview [129, 188, 138, 57, 58, 32, 100, 110]. Intimacy [183]. Introduction [1, 78, 125, 141]. Invented [139]. invention [181]. Invisible [131]. Iperbole} [72]. IPO [137]. IPv6 [40]. ISBN [189, 70]. ISBN-978-1-4331-4710-4 [70]. issues [108]. Italian [72].

J [174]. Jackie [169]. January [179]. Japan [146]. JARING [143]. John [172, 28]. Joseph [49, 169]. journalistic [172]. Julien [60]. justice [147]. justice-seeking [147].

Kaminska [129]. Katrin [189]. Kevin [60]. key [190]. Kilnam [151]. Kleinrock [44]. Know [170]. knowledge [41, 178]. Koerner [169]. Korea [145]. Korean [144].

labor [166]. labour [180]. Lang [70]. language [4, 56]. Lann [69]. large [83]. large-scale [83]. Larry [68, 54]. law [126, 104]. Laws [183]. Laying [115]. LCHC [31]. led [68]. Lee [138]. legacies [96]. Legacy [94, 98, 65, 99]. Legitimacy [87, 190]. Leonard [44]. lessons [37]. Let [191]. letter [53]. library [181]. life [161, 71]. lifetime [166]. lineage [74]. list [39]. Lithuania [193]. Little [127]. local [97]. Long [161]. Long-Gone [161]. looking [39]. lost [96]. Love [183]. Lovink [58]. Lund [50].

M5S [42]. MA [169, 89]. made [77].

magazine [118]. magic [150]. Mailing [39]. mailinglist [81]. Mailland [60]. mainstream [184, 18]. mainstreamisation [187]. Making [172, 179, 43]. Malaysia [143]. man [151]. Mao [142]. Maoist [142]. MapQuest [163]. marginal [184]. market [103, 109]. markets [106, 75]. material [25]. matter [118]. McKelvey [88]. Media [168, 110, 81, 128, 187, 53]. mediated [147]. mediatised [147]. Meme} [170]. memories [69]. Memory [168]. men [131, 166]. menu [24]. messaging [85]. Metadata [182]. Methods [80, 79, 78, 83]. Michel [69]. mid [72]. mid-1990s [72]. Middle [11]. Milligan [90]. Ministerial [197]. Ministry [46]. Minitel [121, 60]. Minneapolis [88]. Minnesota [88]. misremediation [98]. missing [6]. Mobile [106]. model [11]. modems [66]. Moderation [178, 185]. moderator [138]. modernity [22]. moment [190]. Movement [42, 133, 144, 55, 138]. multidisciplinary [109]. multilingualism [13]. Myspace [157]. Myth [123]. narratives [97]. Natalie [189]. nation [34]. nationalism [144, 192]. nature [24]. negotiating [186]. neoliberal [196]. net [6]. Netherlands [76]. netizens [145, 188, 188]. Network [107, 123, 72, 105, 37, 34, 86]. networked [180, 70]. networking [44, 143]. Networks [132, 194, 104, 120, 105, 96, 151, 52, 35]. neutrality [107, 37]. newsgroup [166]. Nicholas [28]. Niels [90]. nonlinear [25]. nostalgic [157]. Notes [15]. noun [23]. novelty [38]. null [15].

objectives [173]. Olympics [149]. once [40]. ones [171]. Online [183, 125, 41, 146, 198, 46, 21]. open [104]. operationalising [12]. origins [119, 42]. other [105]. Our [61]. Outlining [116]. overseas [149]. own [27].

packet [65]. Paolo [123]. paradoxes [106]. passwords [161]. Past [168, 42]. Paul [89, 27]. people [130]. Performing [80]. periphery [148]. Perl [54]. Perry [172]. perspective [109]. Peter [70]. Peters [34]. photos [43]. pieces [81]. Pierre [57]. pilot [12]. pioneer [27, 44, 68]. place [112]. Platform [60, 84, 156, 159, 73, 167]. platform-as-infrastructure [73]. platforms [155, 154]. PLATO [7]. play [191]. playful [150]. point [121]. policy [121, 107, 37]. political [175, 82, 42]. politics [130, 165, 117, 96]. Polity [189, 174]. poor [151]. popular [144]. Porn [159]. Possessed [88]. Post [137, 193, 19, 142, 22]. post-digital [19]. Post-IPO [137]. post-Mao [142]. post-secular [22]. post-socialist [193]. postmen [65]. potatoes [65]. Power [111, 132, 35]. pp [189, 88]. practices [143]. Practicing [105]. praxis [142]. PRC [142]. pre [7]. pre-Internet [7]. precarity [154]. Precorporation [10]. precursors [146]. prejudice [148]. Press [60, 169, 89, 88, 189, 174]. pride [148]. primary [39, 134]. Prime [197]. privacy [105]. Privatization [183]. process [8, 107, 178]. Production [168, 41]. Professionals [95]. profiles [173, 30]. project [48, 52]. projects [72]. promise [175, 130]. propaganda [194]. proper [23]. protest [35]. protests [149]. protocol [85]. provider [76]. pseudonymity [30]. public [117, 150, 152]. publics [150]. pure [144]. purges [159].

quarantines [165]. queer [159]. quotation [98].

R [174, 165]. r/WatchRedditDie [165]. racial [119]. Radical [142, 138]. RangKoM [143]. Ranking [168]. Re [2, 157]. reactions [56]. Reagle [169]. real [151]. Realm [50]. rebirths [159]. rec.arts.anime [177]. reconstructing [134]. recounting [81]. Recovering [98]. reddit [165, 179]. Reding [110]. related [187]. relay [149]. Relentless [145]. religion [126, 22]. religious [56]. Remembering [188]. Rennie [174]. Renren [161]. representing [171]. research [21, 19, 134, 102, 135]. researcher [113, 176]. Resistance [101, 195]. Responsibility [95]. responsible [95]. Rethinking [96]. retrospective [84]. reveries [33]. Review [21, 50, 49, 60, 91, 189, 28, 169, 89, 70, 88, 123, 90, 174, 33, 35, 22, 34, 119]. reviewers [36, 63, 93, 124, 153]. revisionism [3]. revisiting [149, 162]. Revolution [49, 169]. rewards [137]. Rieder [100]. right [145]. rights [55, 166]. riot [179]. Rise [72, 24, 163]. Rising [48]. risks [137]. Roberts [68]. romanticism [146]. Ronda [188]. Roots [86, 67]. round [79, 109]. round-doc [79, 109]. roundtable [155]. Routing [86]. running [85]. Russian [192].

S [122]. S-curve [122]. sacred [22]. SAGE [90]. Sam [22]. Samizdat [192]. Sandvig [21]. saved [160]. scale [83]. scenes [57]. scholarly [180]. Sciences [49]. screencast [17]. scripts [13]. search [24]. Searching [6]. second [176]. secondary [75]. secrets [21]. secular [22]. seeking [147, 167]. selection [76]. selective [118]. self [180]. self-archiving [180]. Selves [70]. Semantics [82]. semiotic [43]. sense [172]. Serres [67]. servitude [166]. SGML [98]. Shaping [197]. shareholder [137]. sharing [41, 43, 28]. Sheets [186]. short [133]. should [16]. Signal [85]. significance [160]. Siles [70]. Simulmatics [139]. site [29]. situated [38]. Situating [193, 154]. Social [168, 95, 7]. socialist [193]. society [110, 114, 86]. Socio [75, 47]. Socio-technical [75, 47]. Socrate} [72]. software [182]. something [191]. sonic [128]. Sorting [168]. sources [39, 134]. Soviet [34]. Speech [138, 121]. speeches [197]. sphere [150]. Spheres [22]. stake [184]. standardisation [85]. Standardising [85]. standards [98]. Star [42]. state [144]. States [70]. STEM [194]. stopped [24]. Stories [169, 142]. story [77]. strategies [105]. streaming [74]. structuralism [66]. structure [9]. structures [105]. struggles [29]. studies [83, 135, 60]. Study [80, 38, 43, 12]. studying [21]. Style [186]. subreddits [178]. Surveillance [111]. swarms [25]. Swartz [196]. switching [65]. system [136]. systems [38, 94].

talk [164]. Talking [53]. Taxing [108]. Team [160]. teargas [35]. Technical [86, 47, 75]. technician [138]. techniques [100]. technocratic [133]. technological [154]. technologies [126, 83, 186, 22]. Technology [175, 56, 43, 54]. Tell [20, 10]. Téramond [32]. text [98]. Thanks [36, 63, 93, 124, 153]. therefore [138]. They’re [162]. Thomas [174]. threads [191]. tie [198]. Tiidenberg [189]. timesharing [38]. Tom [157]. Tomsk [97]. tools [128, 43]. Tor [190, 87]. torch [149]. toxic [185]. Tracing [117, 46, 181]. tracking [83, 135]. tragedy [40]. Trajectories [70, 100]. transformations [75]. Transformed [61]. transforming [102]. transnational [52]. treated [16]. trials [190]. Trolls [187]. Tufekci [35]. Tumblr [160, 189]. turn [196]. turning [121, 51]. turning-point [121]. Twitter [35].

U.S [169]. UK [46]. unclaimed [98].

uncovering [7]. uneasy [34]. unexplored [12]. Union [107]. United [70]. University [88]. Unix [151]. untold [77]. uptake [120]. useful [39]. Usenet [188, 80]. user [136]. usernames [30]. users [106]. Using [80]. utopianism [95].

value [137, 128, 18]. VarageSale [75]. Vegas [44]. VidAngel [126]. Video [61]. videos [43]. view [8]. views [197]. vigilantism [147]. Village [162]. villains [145]. visibility [152]. visions [143]. Viviane [110]. vocabulary [187].

Wall [54]. war [194]. wars [39]. WatchRedditDie [165]. ways [54]. Weaving [87]. Web [91, 98, 84, 187, 181, 59, 48, 158, 92, 157, 83, 160, 90, 170, 102, 122, 46, 54, 76, 186, 18, 200, 134, 167, 17, 51, 87, 76, 62]. Web-based [134]. Weibo [117]. Welcome [60, 121]. Werner [33]. where [2]. who [77]. Wi [174]. Wi-Fi [174]. Wide [200, 62]. Wikipedia [190, 50, 169]. Wilken [174]. winning [44]. Wired [118]. within [135]. Women [41, 77, 131, 177]. women-exclusionary [177]. words [27]. Work [50, 134]. World [61, 95, 131, 197, 200, 62, 33]. write [5].

yak [164]. Yakety [164]. Years [32, 191, 62].

Yelp [162]. York [70]. Yugoslavia [194].

Zen [186]. zeros [171]. Zeynep [35].

References

Brugger:2017:IIH

[1] Niels Brügger, Gerard Goggin, Ian Milligan, and Valérie Schafer. Introduction: Internet histories. Internet Histories, 1(1–2):1–7, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1317128.

Abbate:2017:WWI

[2] Janet Abbate. What and where is the Internet? (re)defining Internet histories. Internet Histories, 1(1–2):8–14, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1305836.

Russell:2017:HRB

[3] Andrew L. Russell. Hagiography, revisionism & blasphemy in Internet histories. Internet Histories, 1 (1–2):15–25, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1298229.

Weber:2017:CL

[4] Marc Weber. A common language. Internet Histories, 1(1–2):26–38, 2017. CODEN ???? ISSN 2470-1483.

URL http://www.tandfonline.com/
doi/full/10.1080/24701475.2017.
1317118.

Turner:2017:CWW

[5] Fred Turner. Can we write a cultural history of the Internet? If so, how? Internet Histories, 1 (1–2):39–46, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1307540.

Driscoll:2017:SMN

[6] Kevin Driscoll and Camille Paloque-Berges. Searching for missing “net histories”. Internet Histories, 1 (1–2):47–59, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1307541.

Jones:2017:PCU

[7] Steve Jones and Guillaume Latzko-Toth. Out from the PLATO cave: uncovering the pre-Internet history of social computing. Internet Histories, 1(1–2):60–69, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1307544.

Braman:2017:IHV

[8] Sandra Braman. Internet Histories: the view from the design process. Internet Histories, 1(1–2):70–78, 2017. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2017.1305716.

Streeter:2017:ISF

[9] Thomas Streeter. The Internet as a structure of feeling: 1992–1996. Internet Histories, 1(1–2):79–89, 2017. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2017.1306963.

Elmer:2017:PWF

[10] Greg Elmer. Precorporation: or what financialisation can tell us about the histories of the Internet. Internet Histories, 1(1–2):90–96, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1308197.

Allagui:2017:IME

[11] Ilhem Allagui. Internet in the Middle East: an asymmetrical model of development. Internet Histories, 1 (1–2):97–105, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1305715.

Reisdorf:2017:UHO

[12] Bianca C. Reisdorf, William H. Dutton, Whisnu Triwibowo, and Michael E. Nelson. The unexplored history of operationalising digital divides: a pilot study. Internet Histories, 1 (1–2):106–118, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1311165.

McLelland:2017:ECM

[13] Mark McLelland. Early challenges to multilingualism on the Internet: the case of Han characterbased scripts. Internet Histories, 1 (1–2):119–128, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1280889.

Wasserman:2017:AHI

[14] Herman Wasserman. African histories of the Internet. Internet Histories, 1(1–2):129–137, 2017. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2017.1308198.

Brunton:2017:NDN

[15] Finn Brunton. Notes from /dev/null. Internet Histories, 1(1–2):138–145, 2017. CODEN ???? ISSN 2470-1483.

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/24701475.2017.
                                        1307059.

Alberts:2017:AAD

[16] Gerard Alberts, Marc Went, and Robert Jansma. Archaeology of the Amsterdam digital city; why digital data are dynamic and should be treated accordingly. Internet Histories, 1 (1–2):146–159, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2017.1309852.

Rogers:2017:DWH

[17] Richard Rogers. Doing Web history with the Internet Archive: screencast documentaries. Internet Histories, 1 (1–2):160–172, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2017.1307542.

Winters:2017:BMD

[18] Jane Winters. Breaking in to the mainstream: demonstrating the value of Internet (and web) histories. Internet Histories, 1(1–2):173–179, 2017. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2017.1305713.

Fortunati:2017:DPD

[19] Leopoldina Fortunati. For a dynamic and post-digital history of the Internet: a research agenda. Internet Histories, 1(1–2):180–187, 2017. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2017.1305714.

Schafer:2017:TUA

[20] Valérie Schafer. Tell us about …. Internet Histories, 1(1–2):188–196, 2017. CODEN ???? ISSN 2470-1483.

URL http://www.tandfonline.com/
doi/full/10.1080/24701475.2017.
1301132.

Acker:2017:BRD

[21] Amelia Acker. Book review: Digital research confidential: the secrets of studying behavior online, by Hargittai, E. and Sandvig, C. Internet Histories, 1(1–2):197–199, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1313928.

Webster:2017:BRT

[22] Peter Webster. Book review: Technologies of religion. Spheres of the sacred in a post-secular modernity, by Sam Han. Internet Histories, 1 (1–2):200–201, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1316535.

Bay:2017:WIC

[23] Morten Bay. What is “Internet”? The case for the proper noun and why it is important. Internet Histories, 1(3):203–218, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1339860.

Kerssens:2017:WSE

[24] Niels Kerssens. When search engines stopped being human: menu interfaces and the rise of the ideological nature of algorithmic search. Internet Histories, 1(3):219–237, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2017.1337666.

Micali:2017:TNM

[25] Alberto Micali. Towards a nonlinear, material history of digital swarms. Internet Histories, 1(3):238–257, 2017. CO-DEN ???? ISSN 2470-1483. URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/24701475.2017.1344051.

LeGallic:2017:IFC

[26] Stéphanie Le Gallic. Infonie: a forced conversion to the Internet (1995–2000). Internet Histories, 1 (3):258–272, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2017.1344417.

Bay:2017:CPP

[27] Morten Bay. Conversations with a pioneer: Paul Baran in his own words. Internet Histories, 1(3):273–282, 2017. CODEN ???? ISSN 2470-1483.

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/24701475.2017.
                                        1345469.

Halavais:2017:BRA

[28] Alexander Halavais. Book review: The age of sharing, by Nicholas A. John. Internet Histories, 1 (3):283–284, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2017.1346223.

Hellegren:2017:HCD

[29] Z. Isadora Hellegren. A history of crypto-discourse: encryption as a site of struggles to define Internet freedom. Internet Histories, 1 (4):285–311, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1387466.

vanderNagel:2017:UPD

[30] Emily van der Nagel. From usernames to profiles: the development of pseudonymity in Internet communication. Internet Histories, 1 (4):312–331, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1389548.

Brown:2017:LEH

[31] Katherine Brown. LCHC: the evolution of a hybrid collaboratory. Internet Histories, 1(4):332–348, 2017. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2017.1384111.

Siles:2017:YIC

[32] Ignacio Siles. 25 years of the Internet in Central America: an interview with Guy de Téramond. Internet Histories, 1(4):349–358, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1383733.

Rouyan:2017:FRB

[33] Anahita Rouyan. Film review: Lo and behold: reveries of the connected world, directed by Werner Herzog. Internet Histories, 1(4):359–360, 2017. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2017.1382827.

West:2017:BRH

[34] Diana Kurkovsky West. Book review: How not to network a nation: the uneasy history of the Soviet Internet, by Benjamin Peters. Internet Histories, 1(4):361–363, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2017.1382828.

Summers:2017:BRT

[35] Ed Summers. Book review: Twitter and teargas: the power and fragility of networks protest, by Zeynep Tufekci. Internet Histories, 1(4):364–367, 2017. CO-DEN ???? ISSN 2470-1483. URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/24701475.2017.1383639.

Anonymous:2017:TR

[36] Anonymous. Thanks to reviewers. Internet Histories, 1(4):368–369, 2017. CO-DEN ???? ISSN 2470-1483. URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/24701475.2017.1394109.

Mailland:2018:BIP

[37] Julien Mailland. Building Internet policy on history: lessons of the forgotten 1981 network neutrality debate. Internet Histories, 2(1–2):1–19, 2018. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2018.1450702.

DePaoli:2018:ECN

[38] Stefano De Paoli. The engineer– criminologist and “the novelty of cybercrime”: a situated genealogical study of timesharing systems. Internet Histories, 2(1–2):20–37, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2018.1448498.

Hocquet:2018:MLA

[39] Alexandre Hocquet and Frédéric Wieber. Mailing list archives as useful primary sources for historians: looking for flame wars. Internet Histories, 2(1–2):38–54, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2018.1456741.

Dourish:2018:OFI

[40] Paul Dourish. The once and future Internet: infrastructural tragedy and ambiguity in the case of IPv6. Internet Histories, 2(1–2):55–74, 2018. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2018.1455311.

Fortunati:2018:WKC

[41] Leopoldina Fortunati. Women’s knowledge co-production and sharing in online communities. Internet Histories, 2(1–2):75–97, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2018.1459132.

Musso:2018:OPD

[42] Marta Musso and Marzia Maccaferri. At the origins of the political discourse of the 5-Star Movement (M5S): Internet, direct democracy and the “future of the past”. Internet Histories, 2(1–2):98–120, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2018.1457295.

Poulsen:2018:BST

[43] Søren Vigild Poulsen. Becoming a semiotic technology – a historical study of Instagram’s tools for making and sharing photos and videos. Internet Histories, 2(1–2):121–139, 2018. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2018.1459350.

Bay:2018:CPL

[44] Morten Bay. Conversation with a pioneer: Leonard Kleinrock on the early days of networking, the ARPANET … and winning in Las Vegas. Internet Histories, 2(1–2):140–152, 2018. CO-DEN ???? ISSN 2470-1483. URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/24701475.2018.1446239.

Winters:2018:E

[45] Jane Winters. Editorial. Internet Histories, 2(1–2):153–155, 2018. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2018.1459080.

Raffal:2018:TOD

[46] Harry Raffal. Tracing the online development of the Ministry of Defence and Armed Forces through the UK web archive. Internet Histories, 2 (1–2):156–178, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2018.1456739.

Ben-David:2018:IAS

[47] Anat Ben-David and Adam Amram. The Internet Archive and the socio-technical construction of historical facts. Internet Histories, 2 (1–2):179–201, 2018. CODEN ????

ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2018.1455412.

Kunze:2018:CCE

[48] Svenja Kunze and Brendan Power. Capturing commemoration: the 1916 Easter Rising web archive project. Internet Histories, 2(1–2):202–217, 2018. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2018.1446238.

Cowls:2017:BRI

[49] Josh Cowls and Evan Higgins. Book review: The Internet Revolution in the Sciences and Humanities, by Alan G. Gross and Joseph E. Harmon. Internet Histories, 2(1–2):218–220, 2017. CO-DEN ???? ISSN 2470-1483. URL http:

//www.tandfonline.com/doi/full/
10.1080/24701475.2017.1419841.

Chakraborty:2017:RWW

[50] Anwesha Chakraborty. Review of Wikipedia, Work and Capitalism: A Realm of Freedom?, by Arwid Lund. Internet Histories, 2(1–2):221–223, 2017. CODEN ???? ISSN 2470-1483.

URL http://www.tandfonline.com/
doi/full/10.1080/24701475.2017.
1415279.

Schafer:2018:TDI

[51] Valérie Schafer and Benjamin G. Thierry. The 90s as a turning decade for Internet and the Web. Internet Histories, 2(3–4):225–229, 2018. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2018.1521060.

Siles:2018:ITP

[52] Ignacio Siles. The Internet as a transnational project: connecting Central America through computer networks (1990–1996). Internet Histories, 2(3–4):230–246, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2018.1500793.

Horbinski:2018:TLH

[53] Andrea Horbinski. Talking by letter: the hidden history of female media fans on the 1990s Internet. Internet Histories, 2(3–4):247–263, 2018. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2018.1500794.

Stevenson:2018:HIB

[54] Michael Stevenson. Having it both ways: Larry Wall, Perl and the technology and culture of the early web. Internet Histories, 2(3–4):264–280, 2018. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2018.1495810.

Petin:2018:BDA

[55] Patrick Pétin and Félix Tréguer. Building and defending the alternative Internet: the birth of the digital rights movement in France. Internet Histories, 2(3–4):281–298, 2018. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2018.1521059.

Webster:2018:TER

[56] Peter Webster. Technology, ethics and religious language: early Anglophone Christian reactions to “cyberspace”. Internet Histories, 2(3– 4):299–314, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2018.1468976.

Paloque-Berges:2018:BSI

[57] Camille Paloque-Berges, Victoria Peuvrelle, and Valérie Schafer. Behind the scenes: an interview with Pierre Beyssac. Internet Histories, 2 (3–4):315–327, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2018.1431496.

Schafer:2018:ICD

[58] Valérie Schafer. Interfacing counterculture and digital cultures: an interview with Geert Lovink. Internet Histories, 2(3–4):328–338, 2018. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2018.1526525.

Hockx-Yu:2018:WH

[59] Helen Hockx-Yu. The web as history. Internet Histories, 2(3–4):340–342, 2018. CODEN ???? ISSN 2470-1483.

URL http://www.tandfonline.com/
doi/full/10.1080/24701475.2018.
1509579.

Gebeil:2018:BRJ

[60] Sophie Gebeil. Book review: Julien Mailland & Kevin Driscoll, Minitel Welcome to the Internet, Platform Studies, MIT Press, 2017. Internet Histories, 2(3–4):342–345, 2018. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2018.1496310.

Bonfiglioli:2018:BBH

[61] Rudi Bonfiglioli. Bit by bit: How video games transformed our world by Andrew ervin. Internet Histories, 2 (3–4):346–349, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2018.1449621.

Winters:2018:WHF

[62] Jane Winters. Web 25: histories from the first 25 years of the World Wide Web. Internet Histories, 2 (3–4):350–352, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2018.1439149.

Anonymous:2018:TR

[63] Anonymous. Thanks to reviewers. Internet Histories, 2(3–4):353–354, 2018. CODEN ???? ISSN 2470-1483.

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/24701475.2018.
                                        1535824.

Paloque-Berges:2019:A

[64] Camille Paloque-Bergès and Valérie Schafer. Arpanet (1969–2019). Internet Histories, 3(1):1–14, 2019. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2018.1560921.

Bay:2019:HPP

[65] Morten Bay. Hot potatoes and postmen: how packet switching became ARPANET’s greatest legacy. Internet Histories, 3(1):15–30, 2019. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2018.1544726.

McKelvey:2019:ABD

[66] Fenwick McKelvey and Kevin Driscoll. ARPANET and its boundary devices: modems, IMPs, and the interstructuralism of infrastructures. Internet Histories, 3(1):31–50, 2019. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2018.1548138.

Schafer:2019:BRA

[67] Valérie Schafer, Sarah Cooper, and Camille Paloque-Bergès. Back to the roots of ARPANET and Internet history with Alexandre Serres. Internet Histories, 3(1):51–67, 2019. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2018.1560680.

Bay:2019:CPL

[68] Morten Bay. Conversation with a pioneer: Larry Roberts on how he led the design and construction of the ARPANET. Internet Histories, 3(1):68–80, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2018.1544727.

Paloque-Berges:2019:FMA

[69] Camille Paloque-Bergès and Valérie Schafer. French memories about the ARPANET: a conversation with Michel Élie and Gérard Le Lann. Internet Histories, 3(1):81–97, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2018.1536934.

Locatelli:2019:BRN

[70] Elisabetta Locatelli. Book review: Networked Selves: Trajectories of Blogging in the United States and France, by I. Siles, New York, Peter Lang, 2017. ISBN-978-1-4331-4710-4. Internet Histories, 3(1):98–100, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1565555.

Lauscher:2019:LBH

[71] Anne Lauscher. Life 3.0: being human in the age of artificial intelligence. Internet Histories, 3(1):101–103, 2019. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2019.1565556.

Bory:2019:INH

[72] Paolo Bory. The Italian network hopes: Rise and fall of the Socrate and Iperbole projects in the mid-1990s. Internet Histories, 3(2):105–122, 2019. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2019.1596407.

Helmond:2019:FED

[73] Anne Helmond, David B. Nieborg, and Fernando N. van der Vlist. Facebook’s evolution: development of a platform-as-infrastructure. Internet Histories, 3(2):123–146, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1593667.

Burroughs:2019:CLS

[74] Benjamin Burroughs. A cultural lineage of streaming. Internet Histories, 3(2):147–161, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1576425.

Lingel:2019:STT

[75] Jessa Lingel. Socio-technical transformations in secondary markets: a comparison of craigslist and Varage-Sale. Internet Histories, 3(2):162–179, 2019. CODEN ???? ISSN 2470-1483.

URL http://www.tandfonline.com/
doi/full/10.1080/24701475.2018.
1478267.

Teszelszky:2019:WAN

[76] Kees Teszelszky. Web archaeology in The Netherlands: the selection and harvest of the Dutch web incunables of provider Euronet (1994–2000). Internet Histories, 3(2):180–194, 2019. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2019.1603951.

Hockx-Yu:2019:BBU

[77] Helen Hockx-Yu. Broad band, the untold story of the women who made the Internet. Internet Histories, 3 (2):195–198, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1598078.

Brugger:2019:IIH

[78] Niels Brügger and Ian Milligan. Introduction: Internet histories and computational methods. Internet Histories, 3(3–4):199–201, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1675405.

Brugger:2019:IHC

[79] Niels Brügger, Ian Milligan, Anat Ben-David, Sophie Gebeil, Federico Nanni, Richard Rogers, William J.

Turkel, Matthew S. Weber, and Peter Webster. Internet histories and computational methods: a “rounddoc” discussion. Internet Histories, 3(3–4):202–222, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1639352.

Dame-Griff:2019:HPE

[80] Avery Dame-Griff. Herding the ‘performing elephants:’ using computational methods to study Usenet. Internet Histories, 3(3–4):223–244, 2019. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2019.1652456.

Dieter:2019:CPR

[81] Michael Dieter, David Gauthier, and Marc Tuters. Conversation pieces: On recounting new media art mailinglist cultures. Internet Histories, 3 (3–4):245–274, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1674580.

Hosl:2019:SIP

[82] Maximilian Hösl. Semantics of the Internet: a political history. Internet Histories, 3(3–4):275–292, 2019. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2019.1656921.

Nielsen:2019:ECM

[83] Janne Nielsen. Experimenting with computational methods for large-scale studies of tracking technologies in web archives. Internet Histories, 3 (3–4):293–315, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1671074.

Ben-David:2019:FCP

[84] Anat Ben-David. 2014 not found: a cross-platform approach to retrospective web archiving. Internet Histories, 3(3–4):316–342, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1654290.

Ermoshina:2019:SRC

[85] Ksenia Ermoshina and Francesca Musiani. “Standardising by running code”: the Signal protocol and de facto standardisation in end-to-end encrypted messaging. Internet Histories, 3(3–4):343–363, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1654697.

Fidler:2019:EIR

[86] Bradley Fidler. The evolution of Internet routing: Technical roots of the Network Society. Internet Histories, 3(3–4):364–387, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1661583.

Schriner:2019:WDW

[87] John Schriner. Weaving the Dark Web: Legitimacy on Freenet, Tor, and I2P. Internet Histories, 3(3– 4):388–390, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1623002.

Musiani:2019:BRF

[88] Francesca Musiani. Book review: Fenwick McKelvey. 2018, Internet Daemons. Digital Communications Possessed, Minneapolis: University of Minnesota Press, 320 pp. Internet Histories, 3(3–4):391–393, 2019. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2019.1611241.

Lecuyer:2019:BRG

[89] Christophe Lécuyer. Book review: GPS, by Paul Ceruzzi, MIT Press, Cambridge, MA., 2018. Internet Histories, 3(3–4):394–395, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1629192.

Pedersen:2019:BRS

[90] Max Odsbjerg Pedersen and Helle Strandgaard Jensen. Book review: The SAGE handbook of web history, edited by Niels Brügger and Ian Milligan. Internet Histories, 3(3–4):396–399, 2019. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2019.1657643.

Gebeil:2019:BRC

[91] Sophie Gebeil. Book review: En construction, la Fabrique française d’Internet et du Web dans les années 1990, Ina, coll. Études & Controverses. Internet Histories, 3(3– 4):400–403, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1672952.

Millward:2019:AWD

[92] Gareth Millward. The archived web: doing history in the digital age. Internet Histories, 3(3– 4):404–406, 2019. CODEN ????

ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1673058.

Anonymous:2019:TR

[93] Anonymous. Thanks to reviewers. Internet Histories, 3(3–4):407–408, 2019. CODEN ???? ISSN 2470-1483.

URL http://www.tandfonline.com/
doi/full/10.1080/24701475.2019.
1677399.

Stevenson:2020:LSI

[94] Michael Stevenson and Anne Helmond. Legacy systems: Internet histories of the abandoned, discontinued and forgotten. Internet Histories, 4(1):1–5, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1725854.

Finn:2020:CWD

[95] Megan Finn and Quinn DuPont. From closed world discourse to digital utopianism: the changing face of responsible computing at Computer Professionals for Social Responsibility (1981–1992). Internet Histories, 4(1):6–31, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1725851.

Kerssens:2020:RLI

[96] Niels Kerssens. Rethinking legacies in Internet history: Euronet, lost (inter)networks, EU politics. Internet Histories, 4(1):32–48, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1701919.

Kolozaridi:2020:NWI

[97] Polina Kolozaridi and Dmitry Muravyov. The narratives we inherit: the local and global in Tomsk’s Internet history. Internet Histories, 4(1):49–65, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2020.1723980.

Ammann:2020:RWU

[98] Rudolf Ammann. Recovering the web’s unclaimed legacy of academic text standards: SGML, HTML, and the misremediation of quotation. Internet Histories, 4(1):66–86, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2020.1725853.

Jones:2020:CLC

[99] Meg Leta Jones. Cookies: a legacy of controversy. Internet Histories, 4(1):87–104, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2020.1725852.

Stevenson:2020:HTA

[100] Michael Stevenson and Anne Helmond. The historical trajectories of algorithmic techniques: an interview with Bernhard Rieder. Internet Histories, 4(1):105–114, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2020.1723345.

Byrne:2020:DFC

[101] Helena Byrne. Digital football cultures: Fandom, identities and resistance. Internet Histories, 4(1):115–116, 2020. CO-DEN ???? ISSN 2470-1483. URL http:

//www.tandfonline.com/doi/full/
10.1080/24701475.2019.1678910.

PhD:2020:HAA

[102] Max Kemman PhD. History in the age of abundance? How the web is transforming historical research. Internet Histories, 4(1):117–119, 2020. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2019.1681715.

Fickers:2020:EIE

[103] Andreas Fickers and Valérie Schafer.

Editorial: the Internet and the EU market. Internet Histories, 4 (2):121–124, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1749808.

Henrich-Franke:2020:ECL

[104] Christian Henrich-Franke. EC competition law and the idea of “open networks” (1950s–1980s). Internet Histories, 4(2):125–141, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1743045.

Jones:2020:PPO

[105] Meg Leta Jones and Kevin Ackermann. Practicing privacy on other networks: network structures, economic arrangements, and identity strategies before cookies. Internet Histories, 4(2):142–160, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1747344.

Goggin:2020:MPE

[106] Gerard Goggin. Mobile paradoxes: European emergence of mobile Internet, users, and markets. Internet Histories, 4(2):161–177, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1741968.

Gadringer:2020:NNE

[107] Stefan Gadringer. Network neutrality in the European Union: a communications policy process analysis. Internet Histories, 4(2):178–195, 2020. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2020.1749807.

Danescu:2020:TIA

[108] Elena Danescu. Taxing intangible assets: issues and challenges for a digital Europe. Internet Histories, 4(2):196–216, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1749806.

Schafer:2020:IEM

[109] Valérie Schafer, Andreas Fickers, David Howarth, Francesca Musiani, Julia Pohle, and Dwayne Winseck. “The Internet and the European market” from a multidisciplinary perspective: a “round-doc” discussion. Internet Histories, 4(2):217–227, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1739360.

Danescu:2020:VRH

[110] Elena Danescu. Viviane Reding on her action in the field of the information society and media (2004–2010). Interview by Elena Danescu. Internet Histories, 4(2):228–241, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1742533.

Sam:2020:ASC

[111] Jillet Sarah Sam. The age of surveillance capitalism: The fight for a human future at the new frontier of power. Internet Histories, 4(2):242–244, 2020. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2019.1706138.

Webster:2020:HPD

[112] Peter Webster. A history of place in the digital age. Internet Histories, 4(2):245, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1731250.

Brugger:2020:IHE

[113] Niels Brügger, Gerard Goggin, Ian Milligan, and Valérie Schafer. Internet Histories early career researcher award. Internet Histories, 4(3):247–251, 2020. CO-DEN ???? ISSN 2470-1483. URL http:

//www.tandfonline.com/doi/full/
10.1080/24701475.2020.1778309.

Haristya:2020:ECS

[114] Sherly Haristya. The efficacy of civil society in global Internet governance. Internet Histories, 4(3):252–270, 2020. CODEN ???? ISSN 2470-1483.

URL http://www.tandfonline.com/
doi/full/10.1080/24701475.2020.
1769892.

Grosse:2020:LFC

[115] Meghan Grosse. Laying the foundation for a commercialized Internet: international Internet governance in the 1990s. Internet Histories, 4 (3):271–286, 2020. CODEN ????

ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1769890.

Nunes:2020:OHC

[116] Raul Nunes. Outlining the history of cyberactivism in Brazil. Internet Histories, 4(3):287–303, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1769893.

Jia:2020:TWC

[117] Lianrui Jia and Xiaofei Han. Tracing Weibo (2009–2019): The commercial dissolution of public communication and changing politics. Internet Histories, 4(3):304–332, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1769894.

Ferrari:2020:BMB

[118] Elisabetta Ferrari. Bodies that matter, bodies that don’t: selective disembodiment in the early Wired magazine (1993–1997). Internet Histories, 4(3):333–348, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1769891.

Mari:2020:REE

[119] Will Mari. A review essay: examining the fraught racial, gendered and class-based origins of the early Internet and its antecedents. Internet Histories, 4(3):349–353, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1758436.

Jones:2020:FCU

[120] Steve Jones and David W. Park. The field of communication’s uptake of computers, networks, and the Internet: 1970–2000. Internet Histories, 4(4):355–372, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1704495.

Bellon:2020:GMW

[121] Anne Bellon. “Goodbye Minitel, welcome to the Internet”. The Hourtin speech as a turning-point for French Internet policy. Internet Histories, 4(4):373–389, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1731783.

Priestley:2020:IWE

[122] Maria Priestley, T. J. Sluckin, and Thanassis Tiropanis. Innovation on the web: the end of the S-curve? Internet Histories, 4(4):390–412, 2020. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2020.1747261.

Musiani:2020:BRP

[123] Francesca Musiani. Book review:

Paolo Bory. 2020, The Internet Myth: From the Internet Imaginary to Network Ideologies. Internet Histories, 4(4):413–416, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1806481.

Anonymous:2020:TR

[124] Anonymous. Thanks to reviewers. Internet Histories, 4(4):417–420, 2020. CO-DEN ???? ISSN 2470-1483. URL http:

//www.tandfonline.com/doi/full/
10.1080/24701475.2020.1840157.

Eriksson:2021:GOC

[125] Maria Eriksson and Guillaume Heuguet. Genealogies of online content identification — an introduction. Internet Histories, 5(1):1–7, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2021.1878649.

Feller:2021:VCF

[126] Gavin Feller and Andrew Ventimiglia. VidAngel: Content filtering technologies, religion, and American copyright law. Internet Histories, 5(1):8–29, 2021. CODEN ???? ISSN 2470-1483.

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/24701475.2020.
                                        1831198.

Justie:2021:LHC

[127] Brian Justie. Little history of CAPTCHA. Internet Histories, 5 (1):30–47, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2020.1831197.

Ernst:2021:MEV

[128] Wolfgang Ernst. The media epistemic value of sonic analytics tools. A commentary. Internet Histories, 5(1):48–56, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2020.1862528.

Eriksson:2021:IAK

[129] Maria Eriksson and Guillaume Heuguet. Interview with Aleksandra Kaminska. Internet Histories, 5(1):57–70, 2021. CODEN ???? ISSN 2470-1483.

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/24701475.2021.
                                        1878650.

Bodker:2021:IPP

[130] Henrik Bødker. An Internet for the people: the politics and promise of craigslist. Internet Histories, 5 (1):71–73, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1827615.

Hockx-Yu:2021:IWD

[131] Helen Hockx-Yu. Invisible women, data bias in a world designed for men. Internet Histories, 5(1):74–78, 2021. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2020.1837580.

Flensburg:2021:NPA

[132] Sofie Flensburg and Signe Sophus Lai. Networks of power. Analysing the evolution of the Danish Internet infrastructure. Internet Histories, 5(2):79–100, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1759010.

Beltramini:2021:ATA

[133] Enrico Beltramini. Against technocratic authoritarianism. A short intellectual history of the cypherpunk movement. Internet Histories, 5 (2):101–118, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1731249. See correc-

tion [140] supplying two additional references.

Hodges:2021:FAE

[134] James A. Hodges. Forensic approaches to evaluating primary sources in Internet history research: reconstructing early Web-based archival work (1989–1996). Internet Histories, 5 (2):119–134, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1784539.

Puschmann:2021:FCA

[135] Cornelius Puschmann and Christian Pentzold. A field comes of age: tracking research on the Internet within communication studies, 1994 to 2018. Internet Histories, 5(2):135–153, 2021. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2020.1749805.

Kushner:2021:IUH

[136] Scott Kushner. The instrumentalised user: human, computer, system. Internet Histories, 5(2):154–170, 2021. CO-DEN ???? ISSN 2470-1483. URL http:

//www.tandfonline.com/doi/full/
10.1080/24701475.2020.1810395.

Bilic:2021:GPI

[137] Paško Bilić and Toni Prug. Google’s post-IPO development: risks, rewards, and shareholder value. Internet Histories, 5(2):171–189, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1864959.

Mommeja:2021:ETR

[138] Julie Momméja. “I am an engineer and therefore a radical”: an interview with Lee Felsenstein, from Free Speech Movement technician to Homebrew Computer Club moderator. Internet Histories, 5(2):190–203, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1917903.

Kotliar:2021:ITH

[139] Dan M. Kotliar. If then: How the Simulmatics Corporation invented the future. Internet Histories, 5 (2):204–206, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1878651.

Anonymous:2021:C

[140] Anonymous. Correction. Internet Histories, 5(2):??, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1860392. See [133].

Goggin:2021:AIH

[141] Gerard Goggin, Haiqing Yu, and Kwang-Suk Lee. Asian Internet histories: an introduction. Internet Histories, 5(3–4):207–213, 2021. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2021.1988239.

Qiu:2021:RPC

[142] Jack Linchuan Qiu and Hongzhe Wang.

Radical praxis of computing in the PRC: forgotten stories from the Maoist to post-Mao era. Internet Histories, 5 (3–4):214–229, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1949817.

Stevens:2021:RJI

[143] Hallam Stevens. From RangKoM and JARING to the Internet: visions and practices of electronic networking in Malaysia, 1983–1996. Internet Histories, 5(3–4):230–247, 2021. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2021.1920126.

Lee:2021:BDS

[144] Kwang-Suk Lee. Between the developmental state and popular nationalism: the pure Hangul movement in the early history of the Korean Internet. Internet Histories, 5 (3–4):248–267, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1951960.

Lee:2021:RVF

[145] Hojeong Lee. Relentless villains or fervent netizens?: The alt-right community in Korea, Ilbe. Internet Histories, 5(3–4):268–286, 2021. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2021.1953756.

Fujioka:2021:DCR

[146] Brett J. Fujioka and Julia R. DeCook. Digital cynical romanticism: Japan’s 2channel and the precursors to online extremist cultures. Internet Histories, 5(3–4):287–303, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1919966.

Huang:2021:MMJ

[147] Qian Huang. The mediated and mediatised justice-seeking: Chinese digital vigilantism from 2006 to 2018. Internet Histories, 5(3–4):304–322, 2021. CO-DEN ???? ISSN 2470-1483. URL http:

//www.tandfonline.com/doi/full/
10.1080/24701475.2021.1919965.

Parray:2021:CPP

[148] M. Imran Parray. Choking the ‘periphery’: pride and prejudice in India’s globalizing Internet imaginary. Internet Histories, 5(3–4):323–340, 2021. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2021.1984732.

Guo:2021:DAA

[149] Shaohua Guo. Discursive activism in the age of BBS: revisiting overseas Chinese protests during the 2008 Beijing Olympics torch relay. Internet Histories, 5(3–4):341–358, 2021. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2021.1951962.

Xie:2021:PSM

[150] Emilie Xie, Maxwell Foxman, and Shuo Xu. From public sphere to magic circle: playful publics on the Chinese Internet. Internet Histories, 5 (3–4):359–375, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1982166.

Paloque-Berges:2021:RPM

[151] Camille Paloque-Bergès. The real “poor man’s Arpanet”? A conversation about Unix networks with Kilnam Chon, godfather of the Asian Internet. Internet Histories, 5(3– 4):376–388, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1982167.

Yu:2021:ECI

[152] Haiqing Yu. The evolution of the Chinese Internet: Creative visibility in the digital public. Internet Histories, 5 (3–4):389–391, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1959213.

Anonymous:2021:TR

[153] Anonymous. Thanks to reviewers. Internet Histories, 5(3–4):392–394, 2021. CODEN ???? ISSN 2470-1483.

URL http://www.tandfonline.com/
doi/full/10.1080/24701475.2021.
2008676.

McCammon:2022:SDD

[154] Muira McCammon and Jessa Lingel. Situating dead-and-dying platforms: technological failure, infrastructural precarity, and digital decline. Internet Histories, 6(1–2):1–13, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2071395.

McCammon:2022:DDP

[155] Muira McCammon, Diami Virgilio, Cody Ogden, Kevin Ackermann, Ethan Zuckerman, Robert Gehl, Saima Akhtar, Sultan Al-Azri, Catherine Knight Steele, Amber M. Hamilton, Anat Ben-David, Sarah Wasserman, Sara Namusoga-Kaale, and Joy Lisi Rankin. Dead-and-dying platforms: a roundtable. Internet Histories, 6(1–2):14–30, 2022. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2022.2071396.

Corry:2021:WDP

[156] Frances Corry. Why does a platform die? Diagnosing platform death at Friendster’s end. Internet Histories, 6(1–2):31–47, 2021. CODEN ????

ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1985360.

Miltner:2021:TUA

[157] Kate M. Miltner and Ysabel Gerrard. “Tom had us all doing frontend web development”: a nostalgic (re)imagining of Myspace. Internet Histories, 6(1–2):48–67, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1985836.

Mikolajewska-Zajac:2022:FDC

[158] Karolina Mikołajewska-Zajac and Attila Márton. The four deaths of Couchsurfing and the changing ecology of the web. Internet Histories, 6(1–2):68–89, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2057751.

Floegel:2021:PBP

[159] Diana Floegel. Porn bans, purges, and rebirths: the biopolitics of platform death in queer fandoms. Internet Histories, 6(1–2):90–112, 2021. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2021.1985833.

Ogden:2021:EIC

[160] Jessica Ogden. “Everything on the Internet can be saved”: Archive Team, Tumblr and the cultural significance of web archiving. Internet Histories, 6(1–2):113–132, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1985835.

Jia:2022:FPL

[161] Lianrui Jia. Forgotten passwords and Long-Gone exes: the life and death of Renren. Internet Histories, 6 (1–2):133–153, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2022.2065153.

Kneese:2022:TDY

[162] Tamara Kneese. “They’re describing Yelp in 1992!”: revisiting the Blacksburg Electronic Village. Internet Histories, 6(1–2):154–171, 2022. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2022.2049069.

Wilken:2022:RFM

[163] Rowan Wilken. The rise and fall of MapQuest. Internet Histories, 6 (1–2):172–190, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2022.2057752.

Montalbano:2021:YYD

[164] Kathryn Montalbano. “Yakety yak: Don’t talk back”: An autopsy of anonymity gone awry. Internet Histories, 6(1–2):191–205, 2021. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2021.2020470.

DeCook:2021:RWP

[165] Julia R. DeCook. r/WatchRedditDie and the politics of reddit’s bans and quarantines. Internet Histories, 6 (1–2):206–222, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2021.1997179.

deConing:2022:LIS

[166] Alexis de Coning. A ‘lifetime of indentured servitude:’ rights, labor, and gender anxieties in a dead men’s rights newsgroup. Internet Histories, 6(1–2):223–236, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2038867.

Mackinnon:2022:DGS

[167] Katie Mackinnon. The death of GeoCities: seeking destruction and platform eulogies in Web archives. Internet Histories, 6(1–2):237–252, 2022. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2022.2051331.

Allmann:2021:SMA

[168] Kira Allmann. Social media and the automatic production of memory: Classification, ranking, and sorting of the past. Internet Histories, 6(1–2):253–256, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1953885.

Hockx-Yu:2021:BRW

[169] Helen Hockx-Yu. Book review:

Wikipedia @ 20, Stories of an Incomplete Revolution, edited by Joseph Reagle and Jackie Koerner, The MIT Press, Cambridge, MA, 2020. U.S. $27.95. Internet Histories, 6 (1–2):257–261, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1932273.

Pettis:2022:KYM

[170] Ben T. Pettis. Know Your Meme and the homogenization of web history. Internet Histories, 6(3):263–279, 2022. CODEN ???? ISSN 2470-1483.

URL http://www.tandfonline.com/
doi/full/10.1080/24701475.2021.
1968657.

Ramati:2022:ABD

[171] Ido Ramati. Aleph-bet, dits-and-dahs, zeros and ones: representing Hebrew in character code. Internet Histories, 6(3):280–297, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1953757.

Buozis:2022:MCS

[172] Michael Buozis. Making common sense of cyberlibertarian ideology: The journalistic consecration of John Perry Barlow. Internet Histories, 6 (3):298–314, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1943994.

Jarvis:2022:CIO

[173] Craig Jarvis. Cypherpunk ideology: objectives, profiles, and influences (1992–1998). Internet Histories, 6(3):315–342, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1935547.

Rikitianskaia:2022:BRT

[174] Maria Rikitianskaia. Book review: Thomas, J., Wilken, R., & Rennie, E. (2021). Wi-Fi. Cambridge: Polity press. Internet Histories, 6 (3):343–344, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2031700.

Aidinoff:2022:PAT

[175] Marc Aidinoff. The promise of access:

Technology, inequality, and the political economy of hope. Internet Histories, 6(3):345–348, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1936780.

Brugger:2022:IHS

[176] Niels Brügger, Gerard Goggin, Ian Milligan, and Valérie Schafer. Internet histories second early career researcher award. Internet Histories, 6(4):349–351, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2146569.

Petit:2022:DFA

[177] Aurélie Petit. “Do female anime fans exist?” The impact of women-exclusionary discourses on rec.arts.anime. Internet Histories, 6(4):352–368, 2022. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2022.2109265.

deAndrade:2022:MAB

[178] Daniela Linkevicius de Andrade and Demival Vasques Filho. Moderation and authority-building process: the dynamics of knowledge creation on history subreddits. Internet Histories, 6(4):369–390, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2110564.

Harel:2022:AMD

[179] Terri Lee Harel. Archives in the making: documenting the January 6 Capitol riot on Reddit. Internet Histories, 6(4):391–411, 2022. CODEN ????

ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2022.2103989.

MacDonald:2022:INS

[180] Corina MacDonald. Imagining networked scholarly communication: selfarchiving, academic labour, and the early Internet. Internet Histories, 6(4):412–431, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2022.2103987.

Hegarty:2022:IAW

[181] Kieran Hegarty. The invention of the archived web: tracing the influence of library frameworks on web archiving infrastructure. Internet Histories, 6(4):432–451, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2022.2103988.

Jansma:2022:MDD

[182] Robert Jansma. Metadata dating the digital city: a software archaeological approach. Internet Histories, 6(4):452–472, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2022.2109285.

Pettis:2022:NLL

[183] Ben T. Pettis. The new laws of love: Online dating and the privatization of intimacy. Internet Histories, 6(4):473–476, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2022.2075156.

Clavert:2023:BMM

[184] Frédéric Clavert and Valérie Schafer. Between marginal and mainstream. Communities and ecosystems at stake. Internet Histories, 7(1):1–5, 2023. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2022.2153399.

Passmann:2023:HCT

[185] Johannes Paßmann, Anne Helmond, and Robert Jansma. From healthy communities to toxic debates: Disqus’ changing ideas about comment moderation. Internet Histories, 7(1):6–26, 2023. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2022.2105123.

Wilson:2023:DNA

[186] Derren Wilson, Saeed-Ul Hassan, Naif Radi Aljohani, Anna Visvizi, and Raheel Nawaz. Demonstrating and negotiating the adoption of web design technologies: Cascading Style Sheets and the CSS Zen Garden. Internet Histories, 7(1):27–46, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2055274.

Gomez-Mejia:2023:MWR

[187] Gustavo Gomez-Mejia. The mainstreamisation of web related vocabulary: “Buzz” and “Trolls” in French media archives. Internet Histories, 7(1):47–75, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2089835.

Miller:2023:RNI

[188] Tristan Miller, Camille Paloque-Bergès, and Avery Dame-Griff. Remembering netizens: an interview with Ronda Hauben, co-author of Netizens: on the history and impact of Usenet and the Internet (1997). Internet Histories, 7(1):76–98, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2022.2123120.

Gekker:2023:BRT

[189] Alex Gekker. Book review: Tumblr, By Katrin Tiidenberg, Natalie Ann Hendry, and Crystal Abidin. Polity press: 2021, 240 pp., ISBN 978-1-5095-4108-9. Internet Histories, 7(1):99–103, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2022.2082060.

Gorissen:2023:WWM

[190] Sebastiaan Gorissen and Robert W. Gehl. When Wikipedia met Tor: trials of legitimacy at a key moment in Internet history. Internet Histories, 7(2):105–121, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2021.2015967.

McKitrick:2023:LPS

[191] Brian McKitrick, Martin Gibbs, Melissa J. Rogerson, Bjørn Nansen, and Charlotte Pierce. Let’s play something awful: a historical analysis of 14 years of threads. Internet Histories, 7(2):122–140, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2022.2084823.

Mikhailov:2023:EIS

[192] Dmitry Mikhailov, Nikolay Ternov, and Ivan Bobrov. Emigration to the Internet: “Samizdat” and the genesis of contemporary Russian nationalism. Internet Histories, 7(2):141–160, 2023. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2022.2137756.

Bareikyte:2023:SII

[193] Miglė Bareikytė. Situating the Internet as infrastructure: the case of postsocialist Lithuania. Internet Histories, 7(2):161–178, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.2000246.

Brautovic:2023:FPW

[194] Mato Brautović. The first propaganda war through computer networks: STEM academia and the breakup of Yugoslavia. Internet Histories, 7 (2):179–198, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1992992.

Kurzmeier:2023:RCD

[195] Michael Kurzmeier. Resistance to the current: the dialectics of hacking. Internet Histories, 7(2):199–201, 2023. CO-DEN ???? ISSN 2470-1483. URL http:

//www.tandfonline.com/doi/full/
10.1080/24701475.2023.2188765.

Buozis:2023:IFF

[196] Michael Buozis. Internet freedom, free culture, and free information: Aaron Swartz and cyberlibertarianism’s neoliberal turn. Internet Histories, 7 (3):203–219, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2023.2235131.

Tandon:2023:SDW

[197] Lasya Tandon. Shaping the digital world: views on the Internet in India in Prime Ministerial speeches (1998–2019). Internet Histories, 7 (3):220–236, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2023.2232212.

John:2023:HFO

[198] Nicholas John and Dekel Katz. A history of features for online tie breaking, 1997–2021. Internet Histories, 7(3):237–253, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2023.2214466.

Nabben:2023:CGI

[199] Kelsie Nabben. Cryptoeconomics as governance: an intellectual history from “Crypto Anarchy” to “Cryptoeconomics”. Internet Histories, 7 (3):254–276, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2023.2183643.

Fomasi:2023:GAB

[200] Martin Fomasi, Deborah Barcella, Eleonora Benecchi, and Gabriele Balbi. Genealogy of an archive. The birth, construction, and development of the World Wide Web collection at CERN. Internet Histories, 7(3):277–294, 2023. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2023.2238254.